FOURIER TRANSFORM MICROWAVE SPECTROSCOPIC AND AB INITIO STUDIES OF THE Kr-CH_{4} AND Ar-CH_{4} VAN DER WAALS COMPLEXES

\underline{YAQIAN LIU} and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB, Canada, T6G 2G2.

Rotational spectra of various isotopomers of the Kr-CH_{4} van der Waals complex were recorded using a pulsed molecular beam Fourier transform microwave spectrometer. Transitions corresponding to two excited internal rotor states, namely the j =1, K =0 and j =2, K =1 states, were measured and assigned, in addition to the previously reported ground state transitions.^{a} The j =1, K =0 state was found to be Coriolis perturbed by a nearby j =1, K =1 state. The spectra of the isotopomers containing ^{83}Kr showed nuclear quadrupole hyperfine structure. Rotational, centrifugal distortion, and quadrupole coupling constants were determined and were used to obtain structural parameters. Ab initio potential energy surfaces of Kr-CH_{4} and Ar-CH_{4} were constructed at the CCSD(T) level of theory with the aug-cc-pVTZ basis set supplemented with bond functions. The dipole moments of the two complexes were also computed at various configurations. The obtained results from the ab initio calculations were compared with those from the experiment.

aY. Liu and W. Jäger, 55th International Symposium on Molecular Spectroscopy, paper RE05 (2000).